# Predictive Inorganic Synthesis based on Machine Learning using Small Data sets: a case study of size-controlled Cu Nanoparticles

### v1

@@ 3.2.1 Ensemble regression
As first generation ensemble model, we consider the complete third-degree polynomial model with LASSO regularization (). This model has an of , and RMSE and MAE of nm and nm, respectively. Despite the very good performance of this first generation model, its main disadvantage is the use of an extensive set of features. With our data set consisting of only 18 samples, the risk of overfitting is significant. In addition, a larger feature set results in a more complex model and, therefore, higher computational cost for inference. With this in mind, the following generations are aimed at reducing the model size drastically to improve generalization and reduce computational cost.
In the second generation, the complexity of the model is strongly reduced by using only the seven features (, , T, , , , ) selected based on the ensemble importance analysis. The ensemble average shows no significant loss in performance compared to the first generation (cf., Table 2), despite the significant reduction in the number of features used. In addition, the training and OOB errors of the ensemble show that the second generation is more generalizable as their difference becomes smaller. The training error of the second generation increases slightly (MAE of nm compared to nm), while the OOB error decreases drastically (MAE goes from nm to nm). The smaller difference between training and OOB error indicates that the model is overfitting less and generalizes better.
In the third generation model, the feature set is expanded with four additional features, aimed at the investigation of the influence of non-polynomial features (based on domain knowledge) on the model performance.
@@
This more complex feature set, gives rise to a slight decline in the performance of the model due to stronger overfitting, leading to poorer generalization. Given the small difference in performance between this and the previous generation, a comprehensive analysis of the ensemble and dominance importance[Azen:PsyMeth2003, Ceder:ChemMater2022] of the features is carried out. Based on this analysis features are selected, aimed at a good balance between sufficient complexity for good predictive quality and a reduced size while retaining good generalization, which resulted in our fourth generation model.
The dominance importance and ensemble importance are considered in an iterative fashion, as done in previous work of one of the authors.[vanpoucke:PolymerInt2022, VerdingP:AdvMaterTech2025] An exhaustive set of sub-models are trained, with one, two or three features (from the full feature list of the third generation model) systematically disabled. This allows us to determine the influence of each individual feature on the model’s performance. Five features are found to have a low importance (%) on average over all models. Therefore, only the six features with an average ensemble importance % are kept for the fourth generation (cf. SI Table IX).
@@
The predicted particle sizes resulting from our fourth generation ensemble regression model are shown in figure 3. In this model, each of the six features has an ensemble importance of % (LASSO regularization, ). The quality of the model is only very slightly below the 19-feature model, showing a significant improvement with respect to overfitting (cf. Table 2). Because the size of the model has been greatly reduced, the MAE for the OOB predictions decreases from nm for the third generation to nm for the fourth generation, a marked improvement. Since the average of the ensemble model is equivalent to a single model instance with averaged coefficients [vanpoucke:JournalApplPhys2020], our fourth generation model can be formulated as follows:
In Equation 1, all feature values are standardized prior to training, meaning the coefficients apply to standardized inputs (cf. SI Table X). When standardized experimental feature values are inserted into Equation 1, the model directly returns the predicted particle size in nm.
@@ 3.2.2 Statistical approach (DoE)
Since the ensemble regression models are built on random subsets of the entire data set, it can be expected that they will generalize better than the classical regression model that was trained on the entire data set. Because individual data points have a large impact in small datasets, models trained on the full dataset are highly sensitive to noise and outliers. Training multiple model instances on random subsets and averaging their predictions reduces this sensitivity, yielding more accurate and robust models.[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022] On the other hand, the classical regression model consists of only two terms containing the precursor concentration, compared to the six features in the ensemble regression model, which results in a simpler model and therefore lower computational cost. Comparison of the ML and DoE models indicates that Cu concentration is an important factor, as both contain it as a polynomial term. In addition, the influence of time and temperature are of relevance as well (as known from experiment) which is captured by the ML models only.
Since the DoE model only retained two features, we attempted to further reduce the fourth generation model by removal of additional features, based on dominance importance analyses (cf. SI Table XI). First was removed, resulting in a slight performance reduction. dropped from to and the average MAE increased from nm to nm. On the OOB data, both models have almost the same MAE, so despite the slightly lower performance there is no increase in generalization. Next, also was removed, as it was assumed to be covered by the feature. Now the overall performance further dropped with an of and an average MAE of nm. In this case, generalization does increase significantly with a training MAE of nm and an OOB MAE of nm. Finally, also was removed from the feature list, to arrive at a three feature model. Again, the generalization of the model increases with a training MAE of nm and an OOB MAE of nm, while maintaining similar average model performance with an of and an average MAE of nm.
@@ 3.2.3 Synthesis prediction: model validation
For three different target sizes ( nm, nm and nm) Equation 2 is used to suggest two possible syntheses routes for each (cf. SI Table XII). The suggested syntheses are performed using the same experimental procedure used for creating the original data set samples. For two of the targeted particle sizes one of the suggested syntheses routes closely matched the target size within experimental uncertainty (cf. samples & in Table 3). For the largest particles (samples & ) neither suggested set of parameters gave the expected results. It is interesting to note that for the incorrect results, the particles are always too small. The suggested synthesis are also used to predict particle size using the DoE model (Equation 3). This shows that the DoE model exhibits a flatter, less pronounced curvature in the explored parameter space compared to the ML model.
4 Conclusion
@@
Acknowledgment
The computational resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government–department WEWIS.

### v2

@@ 3.2.1 Ensemble regression
As first generation ensemble model, we consider the complete third-degree polynomial model with LASSO regularization (). (cf. SI Figure 15) This model has an of , and RMSE and MAE of nm and nm, respectively. (cf., Table 2) Despite the very good performance of this first generation model, its main disadvantage is the use of an extensive set of features. With our data set consisting of only 18 samples, the risk of overfitting is significant. In addition, a larger feature set results in a more complex model and, therefore, higher computational cost for inference. With this in mind, the following generations are aimed at reducing the model size drastically to improve generalization and reduce computational cost.
In the second generation, the complexity of the model is strongly reduced by using only the seven features (, , T, , , , ) selected based on the ensemble importance analysis. The ensemble average shows no significant loss in performance compared to the first generation (cf., Table 2), despite the significant reduction in the number of features used. In addition, the training and OOB errors of the ensemble show that the second generation is more generalizable as their difference becomes smaller. The training error of the second generation increases slightly (MAE of nm compared to nm), while the OOB error decreases drastically (MAE goes from nm to nm). The smaller difference between training and OOB error indicates that the model is overfitting less and generalizes better. (cf. SI Figure 16)
In the third generation model (cf. SI Figure 17), the feature set is expanded with four additional features, aimed at the investigation of the influence of non-polynomial features (based on domain knowledge) on the model performance.
@@
This more complex feature set, gives rise to a slight decline in the performance of the model due to stronger overfitting, leading to poorer generalization. (cf., Table 2) Given the small difference in performance between this and the previous generation, a comprehensive analysis of the ensemble and dominance importance[3, 22] of the features is carried out. Based on this analysis features are selected, aimed at a good balance between sufficient complexity for good predictive quality and a reduced size while retaining good generalization, which resulted in our fourth generation model.
The dominance importance and ensemble importance are considered in an iterative fashion, as done in previous work of one of the authors.[42, 45] An exhaustive set of sub-models are trained, with one, two or three features (from the full feature list of the third generation model) systematically disabled. This allows us to determine the influence of each individual feature on the model’s performance. Five features are found to have a low importance (%) on average over all models. Therefore, only the six features with an average ensemble importance % are kept for the fourth generation (cf. SI Table X).
@@
The predicted particle sizes resulting from our fourth generation ensemble regression model are shown in figure 3. In this model, each of the six features has an ensemble importance of % (LASSO regularization, ). The quality of the model is only very slightly below the 19-feature model, showing a significant improvement with respect to overfitting (cf. Table 2). Because the size of the model has been greatly reduced, the MAE for the OOB predictions decreases from nm for the third generation to nm for the fourth generation, a marked improvement. Since the average of the ensemble model is equivalent to a single model instance with averaged coefficients [29], our fourth generation model can be formulated as follows:
In Equation 1, all feature values are standardized prior to training, meaning the coefficients apply to standardized inputs (cf. SI Table XI). When standardized experimental feature values are inserted into Equation 1, the model directly returns the predicted particle size in nm.
@@ 3.2.2 Statistical approach (DoE)
Since the ensemble regression models are built on random subsets of the entire data set, it can be expected that they will generalize better than the classical regression model that was trained on the entire data set. Because individual data points have a large impact in small datasets, models trained on the full dataset are highly sensitive to noise and outliers. Training multiple model instances on random subsets and averaging their predictions reduces this sensitivity, yielding more accurate and robust models.[29, 42] On the other hand, the classical regression model consists of only two terms containing the precursor concentration, compared to the six features in the ensemble regression model, which results in a simpler model and therefore lower computational cost. Comparison of the ML and DoE models indicates that Cu concentration is an important factor, as both contain it as a polynomial term. In addition, the influence of time and temperature are of relevance as well (as known from experiment) which is captured by the ML models only.
Since the DoE model only retained two features, we attempted to further reduce the fourth generation model by removal of additional features, based on dominance importance analyses (cf. SI Table XII). First was removed, resulting in a slight performance reduction. dropped from to and the average MAE increased from nm to nm. On the OOB data, both models have almost the same MAE, so despite the slightly lower performance there is no increase in generalization. Next, also was removed, as it was assumed to be covered by the feature. Now the overall performance further dropped with an of and an average MAE of nm. In this case, generalization does increase significantly with a training MAE of nm and an OOB MAE of nm. Finally, also was removed from the feature list, to arrive at a three feature model. Again, the generalization of the model increases with a training MAE of nm and an OOB MAE of nm, while maintaining similar average model performance with an of and an average MAE of nm.
@@ 3.2.3 Synthesis prediction: model validation
For three different target sizes ( nm, nm and nm) Equation 2 is used to suggest two possible syntheses routes for each (cf. SI Table XIII). The suggested syntheses are performed using the same experimental procedure used for creating the original data set samples. For two of the targeted particle sizes one of the suggested syntheses routes closely matched the target size within experimental uncertainty (cf. samples & in Table 3). For the largest particles (samples & ) neither suggested set of parameters gave the expected results. It is interesting to note that for the incorrect results, the particles are always too small. The suggested synthesis are also used to predict particle size using the DoE model (Equation 3). This shows that the DoE model exhibits a flatter, less pronounced curvature in the explored parameter space compared to the ML model.
3.3 Classification models
In addition to the regression model, also a binary classification model is created. The goal of this model is to determine whether the obtained Cu NPs will be large (‘1’) or small (‘0’), based on the synthesis parameters. We compared the performance of complex LLMs, trained using transfer learning to a classic random forest model, trained from scratch. Early experiments with epochs are unsuccessful. The models failed to learn the prompt/completion structure, producing invalid outputs for the test data (e.g., numerical artifacts such as ) instead of valid binary labels (‘0’ or ‘1’), preventing the calculation of meaningful performance metrics. With a higher number of epochs, however, the fine-tuned model consistently generated valid predictions (‘0’ or ‘1’). (cf. SI Figure 12)
For LLM-based classification, there is a clear difference between the performance of GPT-J and LLAMA. Where GPT-J performs slightly worse than random guessing, with an accuracy of and a (Table 4), the use of LLAMA 3.1 results in an increase in accuracy to and (Table 4). Although this is a significant performance improvement, the overall performance is still relatively low. One of the possible reasons for this limited performance is the very limited size of the data set. Small data sets already pose major challenges for classic ML, such as regression and decision trees, but the use of such small data sets is particularly challenging for complex models as LLMs. The classic random forest classifier has an overall performance between the two LLMs, with an accuracy of and a (Table 4). In this case, the performance is therefore slightly better than random guessing.
The relatively limited difference in performance between LLMs and classic random forests raises questions about the usefulness of complex neural network based architectures in the context of limited lab-scale studies. It is clear that these require a much larger data set to achieve proper performance.[41] When the data set is limited, as in the current work, there is hardly any difference in their performance compared to random forests. Considering that LLMs are much more complex than classical models makes them less efficient with regard to the computational cost and thus energy requirements.
| Accuracy | ||
| GPT-J | ||
| LLAMA 3.1 | ||
| Random forest |
4 Conclusions
@@
We additionally explored the applicability of LLMs for synthesis prediction in this small-data regime. While LLMs offer conceptual advantages, such as operating directly on recipe-style inputs without explicit feature engineering, our results indicate that they fail to identify reliable relationships between synthesis parameters and synthesis outcome from limited experimental data. While larger data sets might improve LLM performance, generating such data would conflict with the objective of data-efficient synthesis optimization in this current lab-scale experimental setting. These findings indicate that, for lab-scale NP synthesis with data sets typical in experimental materials research, carefully curated small data combined with interpretable classical ML models provides a more reliable approach than deploying more complex LLMs.
@@
Author contributions
Brent Motmans: Data curation, Formal analysis, Investigation, Software, Writing–original draft, Writing–review & editing. Digvijay Ghogare: Supervision, Writing–review & editing. Thijs G. I. van Wijk: Supervision, Writing–review & editing. Joren Van Herck: Formal analysis, Software, Writing–review & editing. Pieter De Meyer: Formal analysis, Software, Writing–review & editing. Berend Smit: Supervision, Writing–review & editing. An Hardy: Conceptualization, Supervision, Writing–review & editing. Danny E. P. Vanpoucke: Conceptualization, Supervision, Writing–review & editing.
Conflicts of interest
There are no conflicts to declare.
Data availability
All data supporting this article are included in the manuscript and Supplementary Information.
Acknowledgements
The authors acknowledge Sander Stulens for performing the selected area electron diffraction (SAED) measurements. This work was financially supported by the Special Research Fund (BOF) via the BOF project, Belgium R15949. The computational resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government–department WEWIS.